OpenStudy (desiraem):

Fan and medal. A sandbag was thrown downward from a building. The function f(t) = -16t2 - 64t + 512 shows the height f(t), in feet, of the sandbag after t seconds. Part A: Factor the function f(t) and use the factors to interpret the meaning of the x-intercept of the function. (4 points) Part B: Complete the square of the expression for f(x) to determine the vertex of the graph of f(x). Would this be a maximum or minimum on the graph? (4 points) Part C: Use your answer in part B to determine the axis of symmetry for f(x)? (2 points)

Vocaloid (vocaloid):

well, first off, do you understand how to factor a quadratic expression?

OpenStudy (desiraem):

No honestly.

Vocaloid (vocaloid):

Hmm, that makes things a bit more difficult...I can try to show you how to do it

Vocaloid (vocaloid):

-16t2 - 64t + 512 = -16(t^2+4t-32), are you with me so far?

OpenStudy (desiraem):

no?

Vocaloid (vocaloid):

all I did was "take out" -16 from each term and rewrote the expression

Vocaloid (vocaloid):

do you agree that 4 + 2 = 2(2 + 1)?

OpenStudy (desiraem):

yes.

Vocaloid (vocaloid):

well, I did the same thing with -16t2 - 64t + 512 -16t2 - 64t + 512 = -16(t^2+4t-32)

Vocaloid (vocaloid):

if you multiply -16(t^2+4t-32) out, you'll get -16t2 - 64t + 512

OpenStudy (desiraem):

okay i got it.

Vocaloid (vocaloid):

great! Let's keep going: t^2+4t-32 = (t+8)(t-4)

OpenStudy (desiraem):

okay

OpenStudy (desiraem):

it's fine, but thanks anyway.